The life expectancy at birth y for females born in a certain country is given by the equation y=0.13x+74.6​, where x is the number of years after 1950.

a. Find the life expectancy for a female born in 1990.

b. Find and interpret the slope of the equation.

c. Find and interpret the​ y-intercept of the equation.


1
Expert's answer
2022-02-10T12:59:37-0500

(a)

 

x = 1990 - 1950 = 40

 

Plug this value of x into the given equation and get y.

 

(b)

 

The slope is the coefficient in front of x, which is 0.13, in units of life expectancy per years since 1950.

 

(c)

 

The y-intercept is the constant in the equation, which is 74.6, in units of years. It represents life expectancy at 1950.

a) y = 0.13(40) + 74.6

   y = 79.8 years

 

b) The slope, 0.13, indicates that the life expectancy for females born in a certain country after 1950 increases by 0.13 years as x increases by 1.

 

c) The y-intercept, 74.6, indicates the life expectancy for females born in a certain country in 1950.